# Lubrication-mediated rebounds off fluid baths

### v1

@@ 3.2 Model Results
Viscosity $\mu_{d}$ 9.78×$\times 10^{-4}$ $\mu_{a}$ 1.825×$\times 10^{-4}$ $\mu_{b}$ 9.78×$\times 10^{-4}$ kgm-1s-1
@@
We first present results of the rebound of two representative spheres of different radii and densities rebounding off a deep water bath against a sweep of initial downward velocities. The choice of parameters is displayed in table 1, and were chosen to correspond to the data for low velocity impacts in Galeano-Rios et al. (2021). This will permit the results of the lubrication-mediated model (LM) derived within this paper to be compared to the kinematic match (KM) model Galeano-Rios et al. (2017) and direct numerical simulation of the Navier-Stokes equations (DNS). In that study experiments were also performed using hydrophobically coated solid spheres.
For such studies, the coefficient of restitution is usually taken as the negative ratio of the speeds between the time of impact $t_{\text{imp}}$, taken to be when the south pole of the sphere first crosses $z=0$ (correspondingly the centre of mass crosses $z=R_{0}$), and the time of liftoff $t_{\text{lift}}$, which is when the sphere leaves passes the same height on an upwards trajectory. The time spent below $z=0$ is the contact time $t_{c}=t_{\text{lift}}-t_{\text{imp}}$. In Galeano-Rios et al. (2017) for low impact velocities, the square coefficient of restitution $\alpha^{2}$, which is the ratio of mechanical energies before and after impact (using the $z=0$ crossing as reference height) is used, as it can capture very small rebounds where $\alpha^{2}<0$ where the sphere detaches but it’s centre of mass does not reach the $z=R_{0}$ after rebound. Thus
@@
where $W_{\text{imp}}$ and $W_{\text{detach}}$ are the vertical velocities at the moment of impact $t_{\text{imp}}$ and separation $t_{\text{detach}}$, when $\text{min}(h)>\epsilon$ is first achieved after impact. In Figure 1, $\alpha^{2}$, the maximum penetration depth $\delta=-(Z_{\text{min}}-R_{0})$, and the pressing time $t_{p}=t_{\text{detach}}-t_{\text{imp}}$, are displayed for the LM and KM models and DNS. In Galeano-Rios et al. (2017) $t_{p}$ is favoured over $t_{c}$ to capture the rebounds which don’t return to $z=0$.
@@
System $\alpha$ $\delta$ $t_{c}$
@@
Figure 4: Comparison between the pressure force felt by the droplet, the pressure along $r=0$ during impact, and the pressed area. Note initial (inertial) and final (suction) spikes. During most of the impact the force is proportional to the lubrication layer area.
@@ 4 Conclusions
The lubrication-mediated model herein efficiently resolves certain features of droplet rebounds heretofore unattainable computationally in reduced models. Future work include computational studies including drop deformation and non-axisymmetric impact.

### v2

@@ 3.2 Model Results
Viscosity $\mu_{d}$ 9.78×$\times 10^{-4}$ $\mu_{a}$ 1.825×$\times 10^{-5}$ $\mu_{b}$ 9.78×$\times 10^{-4}$ kgm-1s-1
Surface tension - - - - $\sigma_{b}$ 7.2×$\times 10^{-2}$ kgs-2
@@
We first present results of the rebound of two representative spheres of different radii and densities rebounding off a deep water bath against a sweep of initial downward velocities. The choice of parameters is displayed in table 1, and were chosen to correspond to the data for low velocity impacts in Galeano-Rios et al. (2021). This will permit the results of the lubrication-mediated model (LM) derived within this paper to be compared to the kinematic match (KM) model (Galeano-Rios et al., 2017) and direct numerical simulation of the Navier-Stokes equations (DNS). In that study experiments were also performed using hydrophobically coated solid spheres.
For such studies, the coefficient of restitution is usually taken as the negative ratio of the speeds between the time of impact $t_{\text{imp}}$, taken to be when the south pole of the sphere first crosses $z=0$ (correspondingly the centre of mass crosses $z=R_{0}$), and the time of liftoff $t_{\text{lift}}$, which is when the sphere leaves the same height on an upwards trajectory. The time spent below $z=0$ is the contact time $t_{c}=t_{\text{lift}}-t_{\text{imp}}$. In Galeano-Rios et al. (2017) for low impact velocities, the square coefficient of restitution $\alpha^{2}$, which is the ratio of mechanical energies before and after impact (using the $z=0$ crossing as reference height) is used, as it can capture very small rebounds where $\alpha^{2}<0$ where the sphere detaches but it’s centre of mass does not reach the $z=R_{0}$ after rebound. Thus
@@
where $W_{\text{imp}}$ and $W_{\text{detach}}$ are the vertical velocities at the moment of impact $t_{\text{imp}}$ and separation $t_{\text{detach}}$, when $\text{min}(h)>\varepsilon$ is first achieved after impact. In Figure 1, $\alpha^{2}$, the maximum penetration depth $\delta=-(Z_{\text{min}}-R_{0})$, and the pressing time $t_{p}=t_{\text{detach}}-t_{\text{imp}}$, are displayed for the LM and KM models and DNS. In Galeano-Rios et al. (2017) $t_{p}$ is favoured over $t_{c}$ to capture the rebounds which don’t return to $z=0$.
@@
System $\alpha$ $\delta$ $t_{p}$
@@
Figure 4: Comparison between the pressure force $F^{*}$ felt by the droplet, the pressure $P^{*}$ along $r=0$ during impact, and the pressed area $A_{p}^{*}$. Note initial (inertial) and final (suction) spikes. The consequences of the force on the wider system is well tracked by the lubrication layer area.
@@ 4 Conclusions
The lubrication-mediated model herein efficiently resolves certain features of droplet rebounds heretofore unattainable computationally in reduced models. Future work include computational studies including drop deformation and non-axisymmetric impact, coupled with the continued validation through results compared to the wider literature (e.g results from studies such as Tang et al. (2018); Alventosa et al. (2023) etc).
@@
[…]
[Declaration of Interests]The authors report no conflict of interest.
[…]
[Funding]This study has been supported by Engineering and Physical Sciences Research Council project no. EP/S022945/1 (K.A.P), and P.A.M gratefully acknowledges support through the Leverhulme project RPG-2023-264.
Appendix A
For simplicity, assuming the drop and bath are the same liquid, the equations can be nondimensionalised using length scale $R$ and surface-tension-based pressure, time, and velocity scales $\sigma/R_{0}$, $(\rho_{b}R_{0}^{3}/\sigma)^{1/2}$ and $(\sigma/\rho_{b}R_{0})^{1/2}$ respectively. The resulting system is
$\partial_{t}\Phi_{b}$ $=-p_{a}-\text{Bo}\,\eta_{b}+\Delta_{H}\eta_{b}+2\,\text{Oh}\,\Delta_{H}\Phi_{b},$ (39)
$\partial_{t}\eta_{b}$ $=D_{b}\Phi_{b}+2\,\text{Oh}\,\Delta_{H}\eta_{b}\doteq F_{b},$ (40)
$\partial_{t}\Phi_{d}$ $=-p_{a}-\kappa_{d}-2\,\text{Oh}\,\partial_{r}^{2}\Phi_{d},$ (41)
$\partial_{t}\zeta_{d}$ $=D_{d}\Phi_{d}+2\,\text{Oh}\left(\frac{1}{r^{2}}\Delta_{s}\zeta_{d}-\frac{1}{r^{3}}\Delta_{s}D_{d}^{-1}\zeta_{d}\right)\doteq F_{d},$ (42)
$\ddot{\mathbf{X}}$ $=F=-\int_{\Omega_{L}}p_{a}\;\mathbf{n_{d}}\;\text{d}a-Bo\;\hat{\mathbf{z}},$ (43)
where all variables are dimensionless and the Bond number $Bo=\rho_{b}gR_{0}^{2}/\sigma$ and the Ohnesorge number $Oh=\mu_{b}/\sqrt{\rho_{b}R_{0}\sigma}$ appear. The derivation of these equations assume small surface deformations (resulting in linear wave systems) and smallness of $Oh$. Indeed the system has already been truncated at leading order in $Oh$ when applying stress balances (Milewski et al., 2015). Consequently, we may therefore formally eliminate any terms $Oh^{2}$ arising from further manipulation of the equations. For the lubrication layer, using a layer thickness $\varepsilon$, we obtain
$\mathbf{Q}=-\frac{\varepsilon^{3}}{R_{0}^{3}}\left(\frac{\mu_{a}}{\sqrt{\rho_{b}R_{0}\sigma}}\right)^{-1}\frac{1}{12}h^{3}\nabla_{H}P+\frac{\varepsilon}{R_{0}}\frac{1}{2}h\left(\nabla_{H}\phi_{b}+\mathbf{u}_{d}^{H}\right),$ (45)
hence the balance $\varepsilon/R_{0}\sim(\mu_{a}/\sqrt{\rho_{b}R_{0}\sigma})^{1/3}$, which is a similar law to that found in Moore (2021), $\varepsilon/R_{0}\sim(\mu_{a}/\rho_{b}R_{0}W_{0})^{1/3}$ for inertially-scaled cushioning, where $W_{0}$ is the impact speed. In our case $\varepsilon/R_{0}\approx 0.05$ the value we have used in simulations.